Andrew lent roger 4800 for 7 months at the end of this period roger had to pay andrew an interest of 119 what was the rate of simple interest per annum?
Answer by rfer(16322)

119/4800=0.02479
0.02479/(7/12)=
0.02479*12/7=0.0425
=4.25%